OpenStudy (anonymous):

the shorter leg of a right triangle is 2cm less then the outer leg. find the length of the two legs if the hypotenuse is 10 cm. I thought I had the answer of 6 and 4 but evidently it is incorrect

OpenStudy (anonymous):

Start with the pythagorean theorem: \[a^{2} + b ^{2} = c ^{2}\] where a is the shortest leg, b is the longer leg, and c is the hypotenuse. Substitute the values into the pythagorean theorem and what do you get? Note: a and b are a different variable in terms of each other.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the equation you get?

OpenStudy (anonymous):

I had x^2+(x-2)^@=10^2

OpenStudy (anonymous):

the @ is supposed to be a

OpenStudy (anonymous):

2

OpenStudy (anonymous):

That's right, so you have to use the quadratic formula, right? What were your steps there?

OpenStudy (anonymous):

x^2+x^2-4+4=100

OpenStudy (anonymous):

2x^2-4-96=0

OpenStudy (anonymous):

2(x^2-2x-48)=0

OpenStudy (anonymous):

(x-2)^2 doesn't equal x^2-4.

OpenStudy (anonymous):

You have to expand it out, using FOIL

OpenStudy (anonymous):

how do i do that my brain is numb now

OpenStudy (anonymous):

FOIL stands for First, Outside, Inside, Last. So, when you have (x-2)^2, that means (x-2)(x-2).

OpenStudy (anonymous):

x^2-x-2 2x+4

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

what is the x-2 2x part?

OpenStudy (anonymous):

x times x and x times -2

OpenStudy (anonymous):

and then 2 times x +2x and 2 times 2 equal 4

OpenStudy (anonymous):

that + was supposed to be an =

OpenStudy (anonymous):

\[(x - 2) (x - 2)\] \[x^{2} - 2x - 2x + 4\] \[x^{2} - 4x + 4\] does this make sense?

OpenStudy (anonymous):

so then you have to add that part... to the original equation. then you'll have a quadratic, so put that in the formula and solve.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x^2-4x-96=0

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

2(x^2-2x-48)=0

OpenStudy (anonymous):

2(x+8)(x-6)=0

OpenStudy (anonymous):

how am I doing so far

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

can someone please help me????????????????????

OpenStudy (anonymous):

sorry I'm here. hang on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay, so far we have: \[x^2+(x-2)^2=10^2\] \[x^2+(x-2)(x-2)=10^2\] \[x^2+x^2-4x+4=10^2\] \[2x^2-4x+4-100=0\] \[2x^2-4x-96=0\]

OpenStudy (anonymous):

Does all that make sense so far?

OpenStudy (anonymous):

yes I have that

OpenStudy (anonymous):

then I divided by 2 so it looks like this 2(x^2-2x-48)=0

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

Yeah, you can do that... but it won't really help you very much. You're trying to solve, so the easiest thing to do is just to put that into the quadratic equation.

OpenStudy (anonymous):

ok so how do I do that

OpenStudy (anonymous):

Quadratic Formula is \[\frac{ -b \pm \sqrt{(b^{2})-4(a)(c)} }{ 2(a) }\]

OpenStudy (anonymous):

omg I don't know how to do that

OpenStudy (anonymous):

hahahaha it's cool. It's easy! I'll walk you through it. So the a, b, and c stand for the different coefficients of the equation.

OpenStudy (anonymous):

For this equation: 2x^2−4x−96=0 a=2 (for the x^2 component) b= -4 (for the x) c= -96 (for the x^0 component)

OpenStudy (anonymous):

2 = a, 4 = b, 96 = c Plug them in to the formula... BAM your x is the result.

OpenStudy (anonymous):

my bad forgot the negative signs, lol.

OpenStudy (anonymous):

yea would have been ugly, without the negatives.

OpenStudy (anonymous):

yeah pretty bad. so plug those in and what do you get?

OpenStudy (anonymous):

that is where I am stuck

OpenStudy (anonymous):

do one first with a + and do a second one with a - in front of the square root.

OpenStudy (anonymous):

Lol yeah I wouldn't do that. solve the big ugly part first, and then you just add or subtract it.

OpenStudy (anonymous):

You will get 2 answers, only one of the answers will be positive you have to use that one because you won't have a negative leg on a triangle.

OpenStudy (anonymous):

http://www.mathwarehouse.com/quadratic/quadratic-formula-calculator.php Plug in: A = 2 B = -4 C = -96

OpenStudy (anonymous):

but what is the equation i am suppose to use

OpenStudy (anonymous):

heyyyy that's cheating!

OpenStudy (anonymous):

(also, if you're going to do that... you might as well send him to wolfram alpha)

OpenStudy (anonymous):

sorry, but she doesn't understand by what we are saying. She is probably in Elem. Algebra in college or in high school Algebra. They don't introduce the Quadratic Formula till the end of Elem. Algebra and sometimes the Beginning of Intermediate Algebra in College.

OpenStudy (anonymous):

i just want to know what to do next so I can finish this ........ yes mat1033 college

OpenStudy (anonymous):

Go to that site and plug in the info I said. Will show you the math.

OpenStudy (anonymous):

I'm 55 and trying to get through this...... I knew I should have taken college and or statistics lol

OpenStudy (anonymous):

Well, write it out, what it looks like when you plug in the numbers... and then I'll show you

OpenStudy (anonymous):

idk what I'm plugging into I don't know what to do next. i thought I did: 2(x^2-2x-48)=0 2(x+8)(x-6)

OpenStudy (anonymous):

I ended up with neg 8 and +6 so I thought the answer was 6,4 but the computer said no

OpenStudy (anonymous):

yeah that's not it at all lol. He gave you the quadratic equation... which is: \[(-b \pm \sqrt{b^2-4ac})\div2a\] where A = 2 B = -4 C = -96

OpenStudy (anonymous):

I have never seen or learned that equation that you are showing.......... never mind I will just get it wrong. I need to go to sleep I have class tomorrow thanks anyway